# SelfAI: Building a Self-Training AI System with LLM Agents

### v1

@@ Abstract
Recent work on autonomous scientific discovery has leveraged LLM-based agents to integrate problem specification, experiment planning, and execution into end-to-end systems. However, these frameworks are often confined to narrow application domains, offer limited real-time interaction with researchers, and lack principled mechanisms for determining when to halt exploration, resulting in inefficiencies, reproducibility challenges, and under-utilized human expertise. To address these gaps, we propose SelfAI, a general multi-agent platform that combines a User Agent for translating high-level research objectives into standardized experimental configurations, a Cognitive Agent powered by LLMs with optimal stopping criteria to iteratively refine hyperparameter searches, and an Experiment Manager responsible for orchestrating parallel, fault-tolerant training workflows across heterogeneous hardware while maintaining a structured knowledge base for continuous feedback. We further introduce two novel evaluation metrics, Score and $\text{AUP}_{D}$, to quantify discovery efficiency and search diversity. Across regression, NLP, computer vision, scientific computing, medical imaging, and drug discovery benchmarks, SelfAI consistently achieves strong performance and reduces redundant trials compared to classical Bayesian optimization and LLM-based baselines, while enabling seamless interaction with human researchers.
1 Main
In recent years, large language models (LLMs) [1, 2, 3] have fundamentally reshaped the landscape of AI-driven research. Advances in reasoning [4, 5], multimodal understanding [6, 7], and autonomous tool use [8, 9, 10, 11, 12, 13] have positioned LLMs as central components for accelerating scientific workflows.
Early LLM-based scientific research demonstrated that LLMs can extract actionable scientific knowledge to guide experiments [14, 15, 16, 17] or answer professional questions [18, 19, 20]. As reasoning and tool-interaction capabilities matured, research expanded to cross-stage planning [21, 22, 23], hypothesis and idea generation [24, 25, 26, 27], multimodal knowledge integration [28], and conducting experiments [17, 29]. Meanwhile, LLMs have driven the development of tasks such as reaction prediction [30, 31, 32], material discovery [33, 34, 18, 35], chemical synthesis design and molecular property optimization [36, 29, 37, 30], and biomedicine research [17, 38, 39, 40, 41]. Building on these advances, a new generation of scientific discovery systems has emerged, spanning AI scientific assistants across the full spectrum of scientific workflows [18, 42, 16, 43], automated scientific discovery systems (ASDSs) specialized for code generation and experiment design [36, 44, 29], and increasingly capable research agents that incorporate automated debugging, iterative idea refinement, and scientific writing [26, 45, 46]. Along this evolutionary trajectory, LLM-centered ASDS have begun to integrate inference, execution, and feedback mechanisms into unified closed-loop pipelines, achieving varying degrees of autonomy across scientific domains.
Despite rapid progress, existing ASDS frameworks still exhibit fundamental limitations. Many systems [44, 45, 24, 47] (such as AIRA and Scientist-V2) excel at translating research intent into executable code or experimental procedures, thereby validating the “executability” of ASDS and primarily relying on final success rate (medal rate) or best performance as metrics. More autonomous platforms like MLGym [48] and MLAgentBench [49] enable system-level control, allowing agents to execute end-to-end benchmarks, collect results, and conduct standardized evaluations. Moreover, while recent LLM-driven optimization methods [50, 51, 52, 53] such as LLM4EO showcase that LLMs can infer evolutionary tendencies and synthesize more effective operators, they operate at the level of local modification rules rather than scientific reasoning in trajectories. Overall, these approaches do not explicitly enhance scientific reasoning during experimentation, such as identifying optimal stopping points, discovering efficient trial trajectories, or assessing the structural quality of reasoning.
Building reasoning-centric ASDS therefore requires a shift from execution-oriented autonomy toward cognitive autonomy: systems must not only carry out experiments but also analyze, reflect, and revise their reasoning throughout the experimental process. However, current ASDS systems provide little support for these cognitive-layer behaviors. The absence of mechanisms for trajectory-level reasoning assessment and LLM-driven decision-making limits both efficiency and accuracy, especially in settings with large search spaces, stochastic outcomes, or complex interdependencies among hypotheses, resulting in inefficient exploration, elevated computational costs, and limited real-time adaptability. These limitations highlight a fundamental need for ASDS architectures that incorporate reasoning-driven search, adaptive learning mechanisms, and structural evaluation of scientific trajectories.
Figure 1: SelfAI Framework for Automated Scientific Experimentation. a, Holistic architecture of the multi-agent system, which transforms various experiments in the research process into a structured workflow. b, User intentions, comprising ideas and experiment schemes, are transformed into structured configurations via a predefined prompt. These inputs are processed through successive stages: hypothesis generation, strategic planning, trial execution, and result collection. c, Performance distribution across 11 tasks demonstrates the framework’s ability, when powered by GPT-4o-mini, to prioritize high-performance regions without sacrificing exploration. d, Trial counts for each solver shown in c, accompanied by quantile lines, density distributions, and performance variability across the global and two evaluation regions. Higher values in low-performance regions promote rapid escape, while lower values in high-performance regions enable localized refinement.

### v2

# SelfAI: A self-directed framework for long-horizon scientific discovery

@@ Abstract
Scientific discovery increasingly entails long-horizon exploration of complex hypothesis spaces, yet most existing approaches emphasize final performance while offering limited insight into how scientific exploration unfolds over time, particularly balancing efficiency-diversity trade-offs and supporting reproducible, human-in-the-loop discovery workflows. We introduce SelfAI, a self-directed, multi-agent-enabled discovery system that automates scientific exploration as a strategic, trajectory-driven decision-making process. SelfAI translates high-level research intent into executable experiments, reasons over accumulated experimental trajectories to guide subsequent exploration, and applies adaptive stopping decisions to terminate unproductive search paths within a closed-loop workflow governed by explicit efficiency-diversity trade-offs. Evaluated using real-world experiments spanning domains from machine learning to drug discovery, SelfAI consistently discovers high-quality solutions with substantially fewer redundant trials than classical optimization and recent LLM-based baselines. The proposed methods establish a general framework for organizing long-horizon scientific discovery and adaptive decision-making in complex scientific and engineering systems.
1 Introduction
Scientific discovery across engineering [bradshaw1983studying, angelopoulos2024transforming], physics [lupoiu2025multi, kaiser2025large], biology [lyu2024alphafold2, nguyen2024sequence], and medicine [yakavets2025machine, grisoni2021combining, jiang2022artificial] increasingly relies on artificial intelligence (AI) systems to support or partially automate key stages of the scientific workflow, ranging from data analysis and hypothesis generation to experimental design and execution. As scientific problems grow in scale, dimensionality, and uncertainty, computational methods have become indispensable for exploring complex hypothesis spaces. AI-assisted discovery systems hold the promise of fundamentally changing the way scientific discovery is practiced [xin2025towards].
Most existing AI-assisted discovery systems [aira, yakavets2025machine, erps2021accelerated] adopt paradigms rooted in machine learning, emphasizing predictive modeling on large datasets and evaluating success primarily through final performance metrics. While effective for many applications, these paradigms provide limited support for reasoning about the discovery process itself, including how hypotheses are explored, how experiments are prioritized, and how decisions are made under uncertainty. In particular, they rarely treat scientific discovery as a sequential, long-horizon decision process in which exploration strategies, resource allocation, and stopping decisions must be jointly optimized over time. Unlocking the full potential of AI in discovery, therefore, requires moving beyond performance-centric automation toward frameworks that explicitly support exploration, decision-making, and adaptation throughout the discovery trajectory.
Recent advances in large language models (LLMs) [openai2023gpt, bai2023qwen, guo2025deepseek] have significantly expanded the potential of AI systems in scientific research. Improvements in reasoning, multimodal understanding, and autonomous tool use now enable LLM-based systems to integrate planning, execution, and feedback within unified workflows. As a result, these systems can support key components of scientific discovery, including hypothesis generation, experiment design, and iterative refinement across domains [huang2022towards, ferrag2025llm]. Building on these capabilities, prior work has progressively enhanced individual aspects of the scientific workflow. Early efforts demonstrated that LLMs can extract actionable scientific knowledge to guide experiments [lin2023evolutionary, angelopoulos2024transforming] and answer domain-specific professional questions [alampara2025probing, polak2024extracting, dagdelen2024structured]. Subsequent research extended these capabilities to cross-stage planning [biomni], automated hypothesis generation [wang2025discovery, aira, baek2025researchagent], multimodal knowledge integration [steyaert2023multimodal, gao2025chemical], and automated experimentation across scientific domains. These advances have led to a new generation of scientific discovery systems, ranging from AI research assistants to fully autonomous workflows capable of executing end-to-end benchmarks [king2004functional, mandal2025evaluating, ai_co_scientist].
While these systems demonstrate the feasibility of system-level automation in scientific workflows, most existing approaches focus primarily on improving individual stages of the discovery process or optimizing final outcomes. Even recent LLM-driven optimization methods [liao2025llm4eo, jiang2025agenticsciml, kochnev2025optuna], which infer evolutionary patterns and synthesize update rules, operate largely at the level of execution or local improvement. However, scientific discovery is inherently a sequential and strategic process that unfolds over time. It requires reasoning not only about isolated steps, but about the structure and quality of the entire exploration trajectory. Current systems rarely treat discovery as a trajectory-level decision-making problem. Consequently, they do not explicitly address fundamental challenges such as determining optimal stopping points, designing efficient exploration paths, or evaluating the structural effectiveness of discovery strategies over time.